Let n E N and A e Mnxn(R). Define trace(A) = x1=1 di,i (i. e. the sum of the diagonal entries) and tr : Mnxn(R) + R, A trace(Compute dim(im(tr)) Enter your answer here and dim(ker(tt)) = Enter your answer here
Answer #1

Consider the linear map tr: M. (IR) IR nxn which maps A E Maxn (R) to its trace. The tr : Max (IR) IR is onto For any XER TheThen trace (A) = x +0+.. ito (n-1) times 3 x That is tr (A) = 4 tr: Mnxn OR) R is Onto.

Then

Range of this map = 1R > im (tr) im (tv) - TR dim (im (tv) - dim (K) = 1 Then That is dim (im (tr)) = 1 We have dim (Maxn GR)

Then by rank -nullity theorem,

Rank(tr) + Nullity (tr) = dim(Mnxn (R)

\Rightarrow\dim (im (tr))+\dim (ker (tr))=\dim(M_{n\times n}(\mathbb{R}))

\Rightarrow\dim (ker (tr))=\dim(M_{n\times n}(\mathbb{R}))-\dim (im (tr))

\Rightarrow\dim (ker (tr))=n^2-1

Hence we get,

\dim (ker (tr))=n^2-1 and \dim (im (tr))=1 .

Rank-Nullity Theorem :

Let  \displaystyle V and \displaystyle W are two vector spaces over a same field \mathbb{F} , where \displaystyle V is finite dimensional. Let \displaystyle T:\displaystyle V\to\displaystyle W be a linear transformation. Then

\operatorname {Rank}(T)+\operatorname{Nullity}(T)=\dim(V)

where \operatorname {Rank}(T)=\dim\left(im (\displaystyle{T})\right) and \operatorname {Nullity}(T)=\dim\left(ker (\displaystyle{T})\right) .

Note :

  • For any nonzero real number a, {a} is a basis for \mathbb{R} , therefore \dim\mathbb{R}=1
  • consider the subset  \left\{A_{ij}:1\leq i,j\leq n\right\} of \displaystyle M_{n\times n}(\mathbb{R}) , where A_{ij}\in\displaystyle M_{n\times n}(\mathbb{R}) such that 1 in the (i,j)^{th } entry and all other entries are zero. Then \left\{A_{ij}:1\leq i,j\leq n\right\} is a basis for \displaystyle M_{n\times n}(\mathbb{R}) and hence  \dim\left(\displaystyle M_{n\times n}(\mathbb{R})\right)=n^2.